A crate of mass = 25 kg is pulled with a force = 140 N and direction = 27• on a horizontal floor . Assume friction is negligible. Calculate:
A. The acceleration of the crate
B. The magnitude of the normal force
1
Expert's answer
2018-10-23T11:41:09-0400

The Newton’s second law states

〖ma=F〗_net

F_net=F cos⁡〖27°〗=140 N×0.891=124.7 N

So, the acceleration of the crate

a=(124.7 N)/(25 kg) =5.0 m/s^2

The magnitude of the normal force

N=mg-F sin⁡〖θ=25kg×9.8 m/s^2 -140N×sin⁡〖27°〗 〗=181 N

Answers: 5.0 m/s^2 ,181 N